OpenStudy (anonymous):

how do i solve 8E+05 * 3E+13

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

okay, thank you for the help. can you help me solve this math problem?

OpenStudy (anonymous):

Is your question \[\huge 8^5 \times 3^{13}\]

OpenStudy (anonymous):

yes, we are currently dealing with scientific notation, so that makes sense. how do i solve? i have a few other questions just like it and i'm super confused :/

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Yes I'm here

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

Okay, so what is 8 times 3?

OpenStudy (anonymous):

It should be this right? \[\huge(8 \times 10^5) \times( 3 \times 10^{13})\]

OpenStudy (anonymous):

hmm. im not sure. that looks right too D:

OpenStudy (anonymous):

Lawl how are we supposed to do then? ;) Wait you've said you were in a scientific notation class right now?

OpenStudy (anonymous):

Okay so its the second one

OpenStudy (anonymous):

well, industry and trades. starting with scientifc notation so yes

OpenStudy (anonymous):

Yeah so, first multiply 8 and 3

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Okay, now, as for exponents, you'll have to add them, what is 5 plus 13?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Okay now you get \[\huge 24 \times 10^{18}\] This is not your answer, you'll have to put it in decimal coz there's 'tens', you need 'ones' So when you decrease the number 24 to 2.4, you add 1 to the exponent, so became 19 \[\huge 2.4 \times 10^{19}\] Done

OpenStudy (anonymous):

thank you so much! can i ask for help on one more, a little more challenging?

OpenStudy (anonymous):

Okay :) you can hit me with anything, I'm up till university level :p

OpenStudy (anonymous):

haha must be nice! the next problem is 2.3E-05 / 1.15E-10

OpenStudy (anonymous):

Okay, so now, you'll have to divide 2.3 with 1.15

OpenStudy (anonymous):

so 2

OpenStudy (anonymous):

Then, you've notice that its dividing, so exponents are to subtract, what is (-5)-(-10)?

OpenStudy (anonymous):

yeah i have no idea. haha

OpenStudy (anonymous):

Well, a minus and a minus gives a plus, so -5-(-10) turns into -5+10 and that you get 5

OpenStudy (anonymous):

The whole thing turns into \[\huge 2 \times 10^5\]

OpenStudy (anonymous):

okay. 8th grade is comin back to me now

OpenStudy (anonymous):

This might be useful \[(+)(+)=(+) \\ (+)(-)=(-) \\ (-)(+)=(-) \\ (-)(-)=(+)\]

OpenStudy (anonymous):

yes it is. i remember that now. so will i then just do 20 times itself five times?

OpenStudy (anonymous):

No don't do that, that is \[2^5=2 \times 2 \times 2 \times 2 \times 2=32\] If its \[2 \times 10^5 \] means \[2 \times 10 \times 10\times 10\times 10\times 10=200000\]

OpenStudy (anonymous):

oh my lord. alright thank you so much for all your help. if i have questions again how do i request you? or can i?

OpenStudy (anonymous):

Of course c:

OpenStudy (anonymous):

and you're welcome :)

OpenStudy (anonymous):

Thank you! expect to hear from me soon haha! have a good day

OpenStudy (anonymous):

good luck and have fun :)